Circular No. 7582

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SAX J1711.6–3808 J. J. M. in ’t Zand, Astronomical Institute, Utrecht University, and

Space Research Organization Netherlands (SRON), Utrecht; R. G. Kaptein, BeppoSAX Science Operation Center, Rome, and SRON; and J. Heise, SRON, communicate: “During an observation with Wide Field Camera unit 1 aboard BeppoSAX on Feb. 8.8–11.5 UT, an x-ray transient was detected at α = 17^{h}11^{m}36^{s}, δ = −38^{o}08′.5 (equinox 2000.0; 99-percent confidence error radius 1′.2). The flux varied between 30 and 80 mCrab (2–9 keV) in hourly integrations. No x-ray bursts were detected. We designate the source SAX J1711.6–3808. Follow-up observations are encouraged.”

COMETS C/2001 C3 AND C/2001 C4 (SOHO) Further to IAUC 7573 and 7580, D. Hammer reports his measurements

for additional Kreutz sungrazing comets found by M. Oates on SOHO website images. Both objects were visible in the C2 coronagraph, but only C/2001 C3 (which also developed a long, thin tail but faded rapidly as it entered the C2 field-of-view on Feb. 4.8 UT) was also visible in C3 data. The reduced observations and orbital elements by B. G. Marsden are included on MPEC 2001-C23.

Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 C3 Feb. 3.488  21^{h}38.^{m}9  −19^{o}17′
C/2001 C4  8.204  21 35.4  −16 01

COMET C/1999 T1 (McNAUGHT-HARTLEY) D. K. Lynch, R. W. Russell, and D. Kim, The Aerospace Corporation;

and M. L. Sitko and S. Brafford, University of Cincinnati, report 3- to 14μm spectroscopy of this comet on Jan. 31.62 and Feb. 1.7 UT using BASS at the Infrared Telescope Facility: “The spectrum on the first night showed a silicate emission feature extending ∼ 12 percent above the continuum defined by a blackbody fitted to the 8- and 13-μm points. Two prominent emission features at 10.3 and 11.2 μm appeared above the silicate band, the latter seemingly indicative of crystalline olivine. The 8- to 13-μm color temperature was 260 ± 10 K, ∼ 10 percent above the blackbody radiative equilibrium temperature of 235 K. The magnitude at 10.5 μm was [N] = 3.0 ± 0.1. On the second night, the two prominent emission features were absent, although the silicate emission feature maintained its trapezoidal shape with breaks at 9.5 and 11.1 μm.”

2001 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green